CARA Telecon Minutes Date: June 18, 2001 Time: 1:00 PM Attendance: BU: Cornell: CMU: Northwestern: R.I.T: S.A.O.: T. Stark South Pole: C. Martin, C. Kaminski Chicago: J. Carlstrom (from OVRO), S. Meyer, R. Pernic, R. Landsberg, J. Bausch UCB: W. Holzapfel (from OVRO) UCSB: J. Ruhl FTS: R. Chamberlin Agenda Items News from the Pole: Chris and Charlie: Instruments working quite well. Viper/ACBAR was broken (AZ drive problem) but fixed itself when it got colder. AST/RO observing at 492 GHz. FTS running; using less helium than expected. Had a noise problem which is being worked on. MAPO building had a heat problem which was fixed after vents were changed around. A glycol smell was detected but seems to have gone away. AST/RO - AST/RO group at SAO is moving to a different building where they will all be together in the same area so it is chaos in Cambridge right now. Will be located in Central Engineering bldg. Acbar - AZ drive sticking; working now that temp has dropped. Will continue to monitor. Sticks at only certain settings, so those will be avoided. SPARO - DASI - continuing to plug away and take good data. No known problems at this point. FTS - Working on calibration issues, using less helium than predicted so will continue another week working out systematics. Helium Report: Run-out date predicted to be between July 24 and August 4. Question still remains as to when to decant the 3000 gallon into smaller storage dewars. There will be a major fill in a few days. *****Charlie will send status report then and final transfer date determined. Question of whether liquid nitrogen could be used to cool down dewars before transferring helium into them; machinery may not be working right but if it is then there should be a daily surplus which could be stored up. Not an option to use 500 gallons which will be remaining at bottom of tank; it is vapor, not liquid. *****John Ruhl will recheck figures from Al's spread sheet; those were mid-May statistics: are they still holding? SPUC meeting: Next Monday-Tuesday. The committee listens to presentations by NSF and Raytheon about what they think is going to happen at SP and the committee makes recommendations. Let Tony know of issues: typically space for scientists, power, communications, liquid helium. Long hold-time storage dewars have been asked for repeatedly at these meetings. Discussion of what type of meeting this would be: Raytheon does a snow job and committee doesn't really have input. Tony will have a committe meeting Sunday to plan strategy. Richard Chamberlin wondered what happened to the 5000 gallon dewar that was on station during his 1995 winter-over? It lasted the whole season. Presumably it was leased, the lease was up and it got sent back. Perhaps Raytheon could look into the 5000 gallon dewar. ** ed note -- this was found to be 3000 G dewar ** Propose SPUC discussion on year-round access on a limited basis be arranged. NSF will say too risky, but it should be discussed. Bob P spoke with Jerry Marty; SIP Review meeting not yet scheduled but he sees deep cuts into tasking, including science tasking. We will go ahead with our plans based on last year's population of 22, and work with AMANDA as necessary.